OpenStudy (anonymous):

Use the factor theorem to match the polynomial with its factor. x3 + x2 + 13x +4

OpenStudy (anonymous):

you mean just factor?

OpenStudy (anonymous):

these are the answer choices x - 5 x+1 x -2 no factor listed for this polynomial x+3

OpenStudy (anonymous):

idk thats the question

OpenStudy (cwrw238):

factor theorem states that if (x-a) is af actor of a polynomial f(x) then f(a) = 0

OpenStudy (anonymous):

im confused

OpenStudy (cwrw238):

so if x+3 is a factor , f(-3) = 0: (-3)^3 + (-3)^2+ 13(-3) + 4 = -27 + 9 -39 + 4 not = 0, so x+3 is not a factor

OpenStudy (anonymous):

where did you get the -3 from if it was a +3?

OpenStudy (cwrw238):

doesn't look like any of the list is a factor

OpenStudy (cwrw238):

no factor listed is the correct choice

OpenStudy (anonymous):

so the answer is no factor listed for this poynomial?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

well wat about x^3 - 4x^2 - 4x - 5

OpenStudy (cwrw238):

because f(5), f(-1) and f(2) not = zero

OpenStudy (anonymous):

the same answers go wit that one too

OpenStudy (cwrw238):

ah - that look more promising as there are negatives in it try f(-1) = -1 - 4 + 4 - 5 = 2 - so x+1 is not a factor

OpenStudy (anonymous):

ohk so basically wat ur doing is plugging in the factors and see if they work??

OpenStudy (anonymous):

i tried -3 and my answer was -8 is tht the rite answer?

OpenStudy (cwrw238):

well for example lets see if x - 5 is a factor - to do this we plug in x=5 f(5) = 125 - 4*25 - 4*5 - 5 = 125 - 100 - 20 -5 = 0 bingo! (x - 5) is a factor

OpenStudy (cwrw238):

yes - 8 - so (x+3) is not a factor

OpenStudy (cwrw238):

you've got the idea now

OpenStudy (anonymous):

ohk so im pluggin in the factors where the x's are and if it equals 0 with tht factor plugged in then it is a factor for tht polynomial and if i dnt get 0 then it is not a factor for tht polynomial? is tht rite?

OpenStudy (cwrw238):

yes - but remember for (x-a) plug in a , but for (x+a) you plug in -a

OpenStudy (cwrw238):

so for example, if f(5) = 0 then (x - 5) is a factor

OpenStudy (anonymous):

ohk thanks